# Combining Neural Fields and Deformation Models for Non-Rigid 3D Motion Reconstruction from Partial Data

### v1

@@ 3.1.4 Reconstruction Refinement
Our method aims to achieve high fidelity reconstructions able to capture fine-grained geometric detail. To that aim, the coarse-dense feature volumes $(\mathcal{F}^{c}_{i})_{i\in\{1,...,N\}}$ are refined where the coarse surface locates using a sparse convolutional decoder [18, 32]. During decoding, fine grained features describing the observed surface retained by the encoder are combined. This allows for high fidelity reconstructions. We denote these refined features $(\mathcal{F}_{i}\in\mathbb{R}^{D_{F}\times H_{F}\times W_{F}\times C})_{i\in\{1,...,N\}}$. As training objectives, we use the same losses defined in Eq. 5 and in Eq. 6 after replacing $(\mathcal{F}^{c}_{i}\in\mathbb{R}^{D_{c}\times H_{c}\times W_{c}\times C})_{i\in\{1,...,N\}}$ with $(\mathcal{F}_{i}\in\mathbb{R}^{D_{F}\times H_{F}\times W_{F}\times C})_{i\in\{1,...,N\}}$ :
@@ 3.2 Inter-Frame Deformation Estimation
For temporal consistency, we guide the reconstruction using a patch-wise near-rigid mesh deformation model [11]. We check that the surface underlying the neural field of each frame can be deformed using this model to obtain the underlying surface of its adjacent frames by optimising for this deformation. This translates the fusion and completion made by $F_{\Psi}$ to a near-isometric deformation in 3D.
One challenge is to unify the representations: the completion operates in a volume, while the deformation operates on a surface. To characterise the surface underlying the geometry aligned features $(\mathcal{F}_{i})_{i\in\{1,...,N\}}$, we extract the zero-level set of the neural fields using marching cubes [36]. This gives a mesh $\mathcal{M}_{i}$ for each frame $i$. The deformation search between neural distances defined by $\mathcal{F}_{i}$ and $\mathcal{F}_{j}$ boils down to a deformation search between meshes $\mathcal{M}_{i}$ and $\mathcal{M}_{j}$.
The mesh deformation model we consider decomposes a non-rigid deformation into patch-wise rigid deformations i.e. a rotation matrix and a translation vector. These patch-wise rigid deformations are blended at the vertex level to obtain the desired non-rigid deformation. Each mesh $\mathcal{M}_{i}$ is therefore decomposed into non-overlapping surface patches $(P^{i}_{k})_{1\leq k\leq L}$ along with their centers $C^{i}=(c^{i}_{k}\in\mathbb{R}^{3})_{1\leq k\leq L}$ where $L$ is the number of patches.
To learn our deformation, we use a patch-wise deformation decoder that takes as input associations between patches of $\mathcal{M}_{i}$ and patches of $\mathcal{M}_{j}$ and outputs rotation and translation parameters that achieve the input associations [39]. Therefore, our deformation search acts in two steps: an association estimation step (Sec. 3.2.1) and a deformation estimation step (Sec. 3.2.2). The inter-frame deformation loss $l^{def}$ in Eq. 1 is composed of an association term and a deformation term:
@@ 3.2.1 Inter-Frame Association Search
Given two meshes $\mathcal{M}_{i}$ and $\mathcal{M}_{j}$ representing neural fields encoded by features $\mathcal{F}_{i}$ and $\mathcal{F}_{j}$, we estimate inter-patch associations in the form of association matrices. We first obtain a feature representative of each patch. This is done by trilinearly interpolating the geometry aligned features $\mathcal{F}_{i}$ and $\mathcal{F}_{j}$ at the center of the patches i.e. $C^{i}$ and $C^{j}$ respectively. This gives for meshes $\mathcal{M}_{i}$ and $\mathcal{M}_{j}$, a feature for each of their patches that we denote $\mathcal{F}^{patch}_{i}\in\mathbb{R}^{L\times C}$ and $\mathcal{F}^{patch}_{j}\in\mathbb{R}^{L\times C}$. Following feature similarity based shape matching methods [16, 28, 39], we use the cosine similarity of these features to estimate inter-patch association matrices, as written by the following equation:
$\footnotesize{(\Pi_{i\rightarrow j})_{mn}}:=\frac{e^{s_{mn}}}{\sum\limits_{k=1}^{L}e^{s_{mk}}}$ (9)
$\footnotesize{(\Pi_{j\rightarrow i})_{mn}}:=\frac{e^{s_{nm}}}{\sum\limits_{k=1}^{L}e^{s_{km}}}$ (10)
@@
For efficiency, we only compute associations and deformations between adjacent frames. We leverage two criteria on our association matrices. First, a cycle consistency criterion [39, 19] promoting cycle consistent associations, i.e. every point going through a cycle is mapped back to itself. We enforce length $2$ and length $3$ cycle consistency for each sequence, which ensures consistency for every cycle [41, 19]. Second, we use a self-reconstruction criterion $l^{rec}$ to identify each patch in feature space, in order to avoid many-to-one patch associations [39]. The combination of $l^{cycle}$ and $l^{rec}$ defines $l^{associate}$ in Eq. 8:
@@
where $\lambda_{3},\lambda_{4}\in\mathbb{R}$ are weights for loss terms. Both $l^{cycle}$ and $l^{rec}$ are detailed in the supplementary (Sec. 9.2.1).
@@ 3.2.2 Deformation Search
The association matrices between meshes $\mathcal{M}_{i}$ and $\mathcal{M}_{j}$ induce our desired deformation. It is the one that deforms $\mathcal{M}_{i}$ (resp. $\mathcal{M}_{j}$) to bring the center of it’s patches from ${C}^{i}$ (resp. ${C}^{j}$) to $\Pi_{i\rightarrow j}{C}^{j}$ (resp. $\Pi_{j\rightarrow i}{C}^{i}$).
@@
The associations were obtained from the geometry aligned features $(\mathcal{F}_{i})_{i\in\{1,...,N\}}$ without any manipulation, hence, the geometry aligned features not only define our geometry when queried through $S_{\Sigma}$, but also define the deformations between the reconstructed surfaces.
@@
To learn this deformation, we employ the deformation decoder introduced for static complete 3D shapes in [39]. It consists of a graph convolutional network acting on the patch neighborhoods followed by an MLP. It outputs rotation parameters $(R_{k}\in\mathbb{R}^{6})_{1\leq k\leq L}$ and new center positions $(u_{k}\in\mathbb{R}^{3})_{1\leq k\leq L}$ for every patch of $\mathcal{M}_{i}$. Applying this deformation leads to the deformed shape $\mathcal{M}_{i\rightarrow j}$.
@@
The deformation network trains using three self-supervised criteria. First, the matching loss encourages the deformation network to match the association matrices by producing the deformations they induce. It is implemented by minimizing:
$\footnotesize\begin{split}l^{match}&=\frac{1}{N-1}(\sum\limits_{i=1}^{N-1}\|C^{i\rightarrow i+1}-\Pi_{i\rightarrow i+1}C^{i+1}\|^{2}_{2}\\ &+\sum\limits_{i=2}^{N}\|C^{i\rightarrow i-1}-\Pi_{i\rightarrow i-1}C^{i-1}\|^{2}_{2}),\end{split}$ (12)
where $C^{i\rightarrow i+1}$ and $C^{i\rightarrow i-1}$ are the deformed cluster centers of $\mathcal{M}_{i\rightarrow i+1}$ and $\mathcal{M}_{i\rightarrow i-1}$ respectively.
@@
Second, the rigidity criterion of the deformation model promotes deformations that preserve the continuity of the deformed shapes along the patch borders:
$\footnotesize l^{rigidity}=\frac{1}{N-1}(\sum\limits_{i=1}^{N-1}l_{rig}(\mathcal{M}_{i\rightarrow i+1})+\sum\limits_{i=2}^{N}l_{rig}(\mathcal{M}_{i\rightarrow i-1})),$ (13)
where $l_{rig}$ is the rigidity loss of the deformation model that we detail in the supplementary (Sec. 9.2.2). The rigidity criterion encourages the deformation network to produce deformations that preserve the intrinsic properties of the mesh i.e.  $l^{rigidity}$ implements the near-isometric assumption.
@@
Third, the surface loss ensures that the deformation produced by the deformation network brings us closer to the surface of the target frame, by encouraging the surface of $\mathcal{M}_{i\rightarrow i+1}$ (resp. $\mathcal{M}_{i\rightarrow i-1}$) to lay on the zero level set of the neural field of frame $i+1$ (resp. frame $i-1$). It is implemented as follows:
@@
$\scriptsize\begin{split}l^{surf}&=\frac{1}{N-1}(\sum\limits_{i=1}^{N-1}(\frac{1}{|V(\mathcal{M}_{i\rightarrow i+1})|}\sum\limits_{\begin{subarray}{c}v\in\\ V(\mathcal{M}_{i\rightarrow i+1})\end{subarray}}|S_{\Sigma}(\operatorname{tri}(\mathcal{F}_{i+1},v),v)|)\\ &+\sum\limits_{i=2}^{N}(\frac{1}{|V(\mathcal{M}_{i\rightarrow i-1})|}\sum\limits_{\begin{subarray}{c}v\in\\ V(\mathcal{M}_{i\rightarrow i-1})\end{subarray}}|S_{\Sigma}(\operatorname{tri}(\mathcal{F}_{i-1},v),v)|)),\end{split}$ (14)
@@ 4 Experiments
We conduct a comparative study on 3D motion reconstruction from monocular depth observations where the goal is to obtain reconstructions with complete shape information and dense tracking. We experiment on both clothed and naked human shapes (Sec. 3) and on animal shapes (Sec. 4.2). We assess with ablations the added benefit of our method’s main components (Sec. 4.3). Supplementary material presents additional quantitative and qualitative results (Sec. 7 and Sec. 8) and implementation details (Sec. 9).
Competing Methods We compare with OFlow [42], LPDC [50], CaDeX [29] and Motion2VecSets [12]. Methods able to train without inter-frame correspondences supervision, i.e. OFlow, CaDeX, and Ours, are trained in the unsupervised regime.
Evaluation Datasets We re-train all methods on the Dynamic FAUST (D-FAUST) [8] dataset, consisting of sequences of minimally dressed, aligned and complete human motion sequences. It includes 10 subjects and 129 sequences. We use the train/val/test split introduced in [42]. To evaluate cross-dataset generalisation, we evaluate the models trained on D-FAUST on two other test sets. First, on a subset of 4DHumanOutfit [5] which consists of sequences of clothed human motions captured using a multi-camera platform. Our subset includes 4 subjects and 8 sequences. Second, on a subset of CAPE [37] which consists of sequences of aligned, complete, and clothed human motions. Our subset includes 3 subjects and 12 sequences.
To evaluate generalisation to other shape classes, we re-train and test all methods on DeformingThings4D-Animals (DT4D-A) [32]. It consists of animations of animal shapes including $38$ animal identities and $1227$ animations. We use the train/val/test split introduced in [29].
For all datasets, we synthetically generate monocular depth videos from the mesh sequences. We use TSDF volumes as input for our method and back-projected depth point-clouds ($10k$ points) for the other methods.
Evaluation Metrics We use the evaluation protocol of [42]. To evaluate the reconstruction and completion, we use Intersection over Union (IoU) and Chamfer Distance (CD). To evaluate tracking, we use the $l_{2}$ correspondence metric (Corr): given a reconstructed sequence and its ground truth, it computes the $l_{2}$ distance between the 3D trajectory of a point on the reconstructions and the trajectory of its corresponding point on the ground truth shapes; correspondence is extracted by a nearest neighbour search to the first frame. Similar to [42, 50, 29, 12], every shape is normalised so the maximum edge length of it bounding box is $1$.
@@ 4.1 Human Motion Sequences
D-FAUST Test Set Table 2 presents the quantitative results on the D-FAUST test set. The test set is divided into two folds containing motions and individuals unseen during training, respectively. In terms of reconstruction and completion quality, our method outperforms all competing methods on both folds. In terms of tracking quality, our method is on par with the other unsupervised method CaDeX, while being close to the best method Motion2VecSets. This shows that departing from the template+flow representation allows to preserve more geometric details while keeping a competitive tracking quality.
Fold Unsup. Input Method IoU $\uparrow$ CD $\downarrow$ Corr $\downarrow$
Back Proj. point cloud LPDC [50] 76.04% 0.00928 0.01176
Mot2VecSets [12] 87.87% 0.00410 0.01014
Unseen Motion OFlow [42] 75.20% 0.00993 0.01648
CaDeX [29] 80.80% 0.00738 0.01191
Mono. Depth TSDF Ours 90.78% 0.00323 0.01342
Back Proj. point cloud LPDC [50] 69.13% 0.01024 0.01392
Mot2VecSets [12] 81.19% 0.00522 0.01155
Unseen Individual OFlow [42] 65.59% 0.01193 0.02050
CaDeX [29] 72.08% 0.00892 0.01480
Mono. Depth TSDF Ours 90.30% 0.00290 0.01245
@@
Table 2: Quantitative comparisons of 4D Shape Reconstruction from monocular depth sequences on D-FAUST [8] (all methods are retrained). Best overall in bold, best amongst unsupervised methods underlined.
Subset of 4DHumanOutfit Table 3 presents the quantitative results on a subset of 4DHumanOutfit. Since the ground truth meshes are not registered to a common template (contrary to D-FAUST and CAPE), we use SMPL [35] fittings to evaluate the tracking. Our method outperforms both supervised and unsupervised methods in reconstruction and tracking, demonstrating superior cross-dataset generalisation. Fig. 3 shows an example. Colors are defined on the first reconstruction and transferred using predicted tracking. OFlow, LPDC and CaDeX fail on this example. Motion2VecSets fails to capture observed details and the unconstrained flow representation causes distortions on unobserved surface parts. Conversely, our representation allows for both preservation of observed geometric detail and for minimising distortions of unobserved surface parts.
Unsup. Input Method IoU $\uparrow$ CD $\downarrow$ Corr $\downarrow$
Back Proj. point cloud LPDC [50] 58.86% 0.01897 0.02739
Motion2VecSets [12] 71.64% 0.00997 0.02583
OFlow [42] 56.03% 0.02154 0.03299
CaDeX [29] 61.28% 0.01804 0.02837
Mono. Depth TSDF Ours 83.14% 0.00584 0.02410
Table 3: Quantitative comparisons of 4D Shape Reconstruction from monocular depth sequences on the 4DHumanOutfit [5] test set (all methods are retrained). Best overall in bold, best amongst unsupervised methods underlined.
Subset of CAPE Table 4 presents the quantitative results on a subset of CAPE. Our method outperforms both supervised and unsupervised methods in reconstruction and tracking.
Unsup. Input Method IoU $\uparrow$ CD $\downarrow$ Corr $\downarrow$
Back Proj. point cloud LPDC [50] 55.51% 0.02085 0.04221
Motion2VecSets [12] 71.94% 0.01140 0.03617
OFlow [42] 50.42% 0.02696 0.05351
CaDeX [29] 57.47% 0.02167 0.04228
Mono. Depth TSDF Ours 85.85% 0.00560 0.03084
@@
Table 4: Quantitative comparisons of 4D Shape Reconstruction from monocular depth sequences on the CAPE [37] test set (all methods are retrained). Best overall in bold, best amongst unsupervised methods underlined.

### v2

@@ 3.1.4 Reconstruction Refinement
We aim to achieve high-fidelity reconstructions that capture fine-grained geometric details. To this end, the coarse-dense feature volumes $(\mathcal{F}^{c}_{i})_{i\in\{1,...,N\}}$ are refined at the coarse surface locations using a sparse convolutional decoder [18, 33]. During decoding, fine-grained features describing the observed surface, retained by the encoder, are combined. This enables high-fidelity reconstructions. We denote these refined features $(\mathcal{F}_{i}\in\mathbb{R}^{D_{F}\times H_{F}\times W_{F}\times C})_{i\in\{1,...,N\}}$. As training objectives, we use the same losses as defined in Eq. 5 and in Eq. 6 after replacing $(\mathcal{F}^{c}_{i}\in\mathbb{R}^{D_{c}\times H_{c}\times W_{c}\times C})_{i\in\{1,...,N\}}$ with $(\mathcal{F}_{i}\in\mathbb{R}^{D_{F}\times H_{F}\times W_{F}\times C})_{i\in\{1,...,N\}}$ :
@@ 3.2 Inter-Frame Deformation Estimation
As neural fields allow for changes of both geometry and topology during training, we leverage the constraint of a near-isometric mesh deformation model [11] to promote topological consistency in a data-driven way. We achieve this by checking that the surface underlying the neural field of each frame can be deformed using this model to get the surface underlying its adjacent frames by optimising for this deformation. This translates the fusion and completion made by $F_{\Psi}$ to a near-isometric deformation in 3D.
One challenge is to unify the representations: the completion acts in a volume, while the deformation acts on a surface. To characterise the surface underlying the geometry aligned features $(\mathcal{F}_{i})_{i\in\{1,...,N\}}$, we extract the zero-level set of the neural fields using marching cubes [37]. This gives a mesh $\mathcal{M}_{i}$ for each frame $i$. The deformation search between neural distances defined by $\mathcal{F}_{i}$ and $\mathcal{F}_{j}$ boils down to a deformation search between meshes $\mathcal{M}_{i}$ and $\mathcal{M}_{j}$. Both representations are then linked using training objectives.
The mesh deformation model we consider decomposes a non-rigid deformation into patch-wise rigid deformations, i.e. a rotation matrix and a translation vector. The latter are blended at the vertex level to obtain the non-rigid deformation. Each mesh $\mathcal{M}_{i}$ is thus decomposed into non-overlapping surface patches $(P^{i}_{k})_{1\leq k\leq L}$ with their centers $C^{i}=(c^{i}_{k}\in\mathbb{R}^{3})_{1\leq k\leq L}$ where $L$ is the number of patches.
To learn the deformation, we use a patch-wise deformation decoder that inputs associations between patches of $\mathcal{M}_{i}$ and patches of $\mathcal{M}_{j}$ and outputs rotation and translation parameters that achieve the input associations [40]. So, the deformation search acts in two steps: an association estimation step (Sec. 3.2.1) and a deformation estimation step (Sec. 3.2.2). The inter-frame deformation loss $l^{def}$ in Eq. 1 is composed of an association term and a deformation term:
@@ 3.2.1 Inter-Frame Association Search
Given two meshes $\mathcal{M}_{i}$ and $\mathcal{M}_{j}$ representing neural fields encoded by $\mathcal{F}_{i}$ and $\mathcal{F}_{j}$, we estimate inter-patch associations as association matrices. We first obtain a feature representative of each patch by trilinearly interpolating $\mathcal{F}_{i}$ and $\mathcal{F}_{j}$ at the center of the patches i.e. at $C^{i}$ and $C^{j}$ respectively. This gives for meshes $\mathcal{M}_{i}$ and $\mathcal{M}_{j}$, a feature for each of their patches denoted $\mathcal{F}^{patch}_{i}\in\mathbb{R}^{L\times C}$ and $\mathcal{F}^{patch}_{j}\in\mathbb{R}^{L\times C}$. Following feature similarity based shape matching methods [16, 29, 40], we use the cosine similarity of these features to estimate inter-patch association matrices as follows:
${(\Pi_{i\rightarrow j})_{mn}}:=\frac{e^{s_{mn}}}{\sum\limits_{k=1}^{L}e^{s_{mk}}}$ (9)
${(\Pi_{j\rightarrow i})_{mn}}:=\frac{e^{s_{nm}}}{\sum\limits_{k=1}^{L}e^{s_{km}}}$ (10)
@@
For efficiency, we only compute associations and deformations between adjacent frames. We leverage two criteria on the association matrices. First, a cycle consistency criterion [40, 19] that promotes cycle consistent associations, i.e. every point that goes through a cycle is mapped back to itself. We enforce length $2$ and length $3$ cycle consistency for each sequence, ensuring consistency for every cycle [43, 19]. Second, we use a self-reconstruction criterion $l^{rec}$ to identify each patch in feature space, in order to avoid many-to-one patch associations [40].
The combination of $l^{cycle}$ and $l^{rec}$ defines $l^{associate}$ in Eq. 8:
@@
where $\lambda_{3},\lambda_{4}\in\mathbb{R}$ are weights for loss terms. Both $l^{cycle}$ and $l^{rec}$ are detailed in Sec. 9.2.1.
@@ 3.2.2 Deformation Search
The association matrices between meshes $\mathcal{M}_{i}$ and $\mathcal{M}_{j}$ induce the desired deformation. These matrices deform $\mathcal{M}_{i}$ (resp. $\mathcal{M}_{j}$) by mapping the center of its patches from ${C}^{i}$ (resp. ${C}^{j}$) to $\Pi_{i\rightarrow j}{C}^{j}$ (resp. $\Pi_{j\rightarrow i}{C}^{i}$).
@@
The associations were obtained from the geometry aligned features $(\mathcal{F}_{i})_{i\in\{1,...,N\}}$ without any manipulation, hence, the geometry aligned features not only define the geometry when queried through $S_{\Sigma}$, but also define the deformations between the reconstructed surfaces.
@@
To learn this deformation, we employ a deformation decoder as introduced in [40] for static, complete 3D shapes. It consists of a graph convolutional network acting on the patch neighborhoods followed by an MLP. It outputs rotation parameters $(R_{k}\in\mathbb{R}^{6})_{1\leq k\leq L}$ and new center positions $(u_{k}\in\mathbb{R}^{3})_{1\leq k\leq L}$ for every patch of $\mathcal{M}_{i}$. Applying this deformation leads to the deformed shape $\mathcal{M}_{i\rightarrow j}$.
@@
The deformation network is trained using three self-supervised criteria. First, the matching loss $l^{match}$ encourages the deformation network to match the association matrices by producing the deformations they induce :
$\begin{split}l^{match}&=\frac{1}{N-1}(\sum\limits_{i=1}^{N-1}\|C^{i\rightarrow i+1}-\Pi_{i\rightarrow i+1}C^{i+1}\|^{2}_{2}\\ &+\sum\limits_{i=2}^{N}\|C^{i\rightarrow i-1}-\Pi_{i\rightarrow i-1}C^{i-1}\|^{2}_{2}),\end{split}$ (12)
where $C^{i\rightarrow i+1}$ and $C^{i\rightarrow i-1}$ are the deformed patch centers of $\mathcal{M}_{i\rightarrow i+1}$ and $\mathcal{M}_{i\rightarrow i-1}$ respectively.
@@
Second, the rigidity criterion promotes preserving the continuity of the deformed shapes along the patch borders:
$l^{rigidity}=\frac{1}{N-1}(\sum\limits_{i=1}^{N-1}l_{rig}(\mathcal{M}_{i\rightarrow i+1})+\sum\limits_{i=2}^{N}l_{rig}(\mathcal{M}_{i\rightarrow i-1})),$ (13)
[…]
@@
Third, the surface loss $l^{surf}$ explicitly links the deforming surfaces with the geometry aligned features. It ensures that the deformation network brings the deformed vertices closer to the surface of the target frame, by encouraging the surface of $\mathcal{M}_{i\rightarrow i+1}$ (resp. $\mathcal{M}_{i\rightarrow i-1}$) to lay on the zero level set of the neural field of frame $i+1$ (resp. frame i1i-1):
@@
$l^{surf}=\frac{1}{N-1}(\sum\limits_{i=1}^{N-1}(\frac{1}{|V(\mathcal{M}_{i\rightarrow i+1})|}\sum\limits_{\begin{subarray}{c}v\in\\ V(\mathcal{M}_{i\rightarrow i+1})\end{subarray}}|S_{\Sigma}(\operatorname{tri}(\mathcal{F}_{i+1},v),v)|)$
$+\sum\limits_{i=2}^{N}(\frac{1}{|V(\mathcal{M}_{i\rightarrow i-1})|}\sum\limits_{\begin{subarray}{c}v\in\\ V(\mathcal{M}_{i\rightarrow i-1})\end{subarray}}|S_{\Sigma}(\operatorname{tri}(\mathcal{F}_{i-1},v),v)|)),$ (14)
@@ 4 Experiments
We conduct a comparative study on 3D motion reconstruction from monocular depth, aiming to recover complete shapes with dense tracking. We test on clothed and naked humans (Sec. 3) and animals (Sec. 4.2). Ablations assess the method’s core components (Sec. 4.3). The supplementary presents additional results in Sec. 7 and implementation details in Sec. 9.
Competing Methods We compare with OFlow [44], LPDC [52], CaDeX [30] and Motion2VecSets [12]. Methods able to train without inter-frame correspondence supervision: OFlow, CaDeX, and Ours, are trained unsupervised.
Evaluation Datasets All methods are trained on the Dynamic FAUST (D-FAUST) [8] dataset, consisting of sequences of minimally dressed, aligned and complete human motion sequences. It includes 10 subjects and 129 sequences. We use the train/val/test split introduced in [44]. To evaluate cross-dataset generalisation, we evaluate the models trained on D-FAUST on two other test sets. First, on a subset of 4DHumanOutfit [5] which consists of sequences of real clothed human motions captured using a multi-camera platform. Our subset includes 4 subjects and 8 sequences. Second, on a subset of CAPE [38] which consists of sequences of aligned, complete, and clothed human motions. Our subset includes 3 subjects and 12 sequences.
To evaluate generalisation to other shape classes, all methods are trained on DeformingThings4D-Animals (DT4D-A) [33]. It consists of animations of animal shapes including $38$ animal identities and $1227$ animations. We use the train/val/test split introduced in [30].
Similar to e.g[30, 12], we generate synthetic monocular depth videos from the mesh sequences. We use back-projected depth point-clouds ($10k$ points) for the competing methods and compute TSDFs for our approach; hence, all methods input densely sampled surface information. For fair comparisons, all methods are re-trained as we generate depth videos that differ from those used in the original works and train on denser point clouds. Still, our results are consistent with the ones reported in the original works.
Evaluation Metrics We use the evaluation protocol of [44]. We evaluate the reconstruction and completion using Intersection over Union (IoU) and Chamfer Distance (CD). For the tracking, we use the $l_{2}$ correspondence metric (Corr): given a reconstructed sequence and its ground truth, it computes the $l_{2}$ distance between the 3D trajectory of a point on the reconstructions and the trajectory of its corresponding point on the ground truth shapes; correspondence is extracted by a nearest neighbour search to the first frame. Similar to [44, 52, 30, 12], every shape is normalised so the maximum edge length of it bounding box is $10$.
@@ 4.1 Human Motion Sequences
D-FAUST Tab. 2 shows the quantitative results on the D-FAUST test set. It is split into motions and individuals unseen during training. In terms of reconstruction and completion quality, our method outperforms all competing methods on both folds. In terms of tracking quality, our method is on par with the other unsupervised method CaDeX, while being close to the best method Motion2VecSets. This shows that departing from the template+flow representation and estimating shape and motion in a coupled manner between frames allows to preserve more geometric details while keeping a competitive tracking quality.
Fold Unsup. Method IoU $\uparrow$ CD $\downarrow$ Corr $\downarrow$
LPDC [52] 76.04% 0.0928 0.1176
Mot2VecSets [12] 87.87% 0.0410 0.1014
Unseen
Motion
OFlow [44] 75.20% 0.0993 0.1648
CaDeX [30] 80.80% 0.0738 0.1191
Ours 90.78% 0.0323 0.1342
LPDC [52] 69.13% 0.1024 0.1392
Mot2VecSets [12] 81.19% 0.0522 0.1155
Unseen
Individual
OFlow [44] 65.59% 0.1193 0.2050
CaDeX [30] 72.08% 0.0892 0.1480
Ours 90.30% 0.0290 0.1245
@@
Table 2: Quantitative 4D shape reconstruction results from monocular depth on D-FAUST [8] (all methods are retrained). Best overall in bold, best amongst unsupervised methods underlined.
Unsup. Method IoU $\uparrow$ CD $\downarrow$ Corr $\downarrow$
LPDC [52] 55.51% 0.2085 0.4221
Motion2VecSets [12] 71.94% 0.1140 0.3617
OFlow [44] 50.42% 0.2696 0.5351
CaDeX [30] 57.47% 0.2167 0.4228
Ours 85.85% 0.0560 0.3084
@@
Table 3: Quantitative 4D shape reconstruction results from monocular depth on CAPE [38] (all methods are retrained). Best overall in bold, best amongst unsupervised methods underlined.
4DHumanOutfit and CAPE Tab. 4 shows quantitative results on a subset of 4DHumanOutfit. Since the ground truth meshes are not registered to a common template (contrary to D-FAUST and CAPE), we use SMPL [36] fittings to evaluate the tracking. Our method outperforms both supervised and unsupervised methods in reconstruction and tracking, demonstrating superior cross-dataset generalisation. Fig. 3 shows an example. OFlow, LPDC and CaDeX fail on this example. Motion2VecSets fails to capture observed details and the unconstrained flow representation causes distortions on unobserved surface parts. Conversely, our representation allows for both preservation of observed geometric detail and for minimising distortions of unobserved surface parts. We report similar results for the CAPE subset in Tab. 3.
Unsup. Method IoU $\uparrow$ CD $\downarrow$ Corr $\downarrow$
LPDC [52] 58.86% 0.1897 0.2739
Motion2VecSets [12] 71.64% 0.0997 0.2583
OFlow [44] 56.03% 0.2154 0.3299
CaDeX [30] 61.28% 0.1804 0.2837
Ours 83.14% 0.0584 0.2410
Table 4: Quantitative 4D shape reconstruction results from monocular depth on 4DHumanOutfit [5] (all methods are retrained). Best overall in bold, best amongst unsupervised methods underlined.
Figure 4: Results with depth map estimates from real monocular RGB images. First and tenth frame reconstructions are shown.
Further qualitative experiments use depth estimates from RGB images using Sapiens [27] as input (see Fig. 4). This is a challenging practical scenario, and our approach produces visually pleasing results.